OpenStudy (anonymous):

The sequence is geometric. What is the next term? 12, 24, 48, 96, ... 186 192 288 108

OpenStudy (anonymous):

What do you do to 12 to turn it into 24? What do you do to 24 to turn it into 48? What do you do to 48 to turn it into 96?

OpenStudy (anonymous):

times it by 2?

OpenStudy (anonymous):

???

OpenStudy (anonymous):

Yep

OpenStudy (anonymous):

sorry im confused

OpenStudy (anonymous):

192

OpenStudy (anonymous):

12+12=24 24+24=48 48+48=96 96+96=192

OpenStudy (anonymous):

ohhhhh okay thankyou i was very confused on how to do that part